OpenStudy (anonymous):

An archery target has three concentric regions. The diameters of the regions are in the ratio 1:2:3. Find the ratio of their areas

OpenStudy (anonymous):

1:4:9 as area is proportional to square of the diameters

OpenStudy (anonymous):

But in the book it said the ratio is 1:3:5

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

we need to find areas of circular strip regions. not the whole circles.

OpenStudy (anonymous):

yes u r right ...my bad..

ganeshie8 (ganeshie8):

diameters are in ratio, \(1 \color{red}{:} 2 \color{red}{:} 3\) ratio of area of regions : \(\pi (d/2)^2 \color{red}{:} \pi (2d/2)^2 - \pi (d/2)^2 \color{red}{:} \pi (3d/2)^2 - \pi (2d/2)^2\)

ganeshie8 (ganeshie8):

see if that makes sense

OpenStudy (anonymous):

I don't understand

ganeshie8 (ganeshie8):

ok area of circle = \(\pi (d/2)^2\)

ganeshie8 (ganeshie8):

ok wid that ?

OpenStudy (anonymous):

I got confused about ratio of area of regions

ganeshie8 (ganeshie8):

yea we will get to that :) so u familiar wid circle area formula already

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):